OpenStudy (pottersheep):

Grade 11 Geometric Progressions help PLEASE~!!!

OpenStudy (pottersheep):

Show that if n is even, Sn = 3Rn , where: Sn = 1 + 1/2 + 1/4 .....(1/2)^(n - 1) and Rn = 1 - 1/2 - 1/4 ......+ (-1/2)^(n-1)

OpenStudy (anonymous):

multiply Rn by three and subtract from Sn ...you got 0 for even valus of n then Sn-3Rn=0 Sn=3Rn